# Hallucination of Multimodal Large Language Models: A Survey

### v1

@@ 2.3. Hallucinations in Multimodal Large Language Models
It’s worth noting that some literature may categorize objects counting, objects event, etc., as independent hallucination categories. In this work, we classify them under the attribute category. The definition of hallucination types aligns well with the domain of compositional generalization (Yuksekgonul et al., 2022; Ma et al., 2023) of VLMs, which investigates visio-linguistic generalization and reasoning abilities.
@@ 3.1.1. Quantity
Deep learning models are data-hungry, especially large models like MLLMs. The amount of data plays an important role in building robust and reliable MLLMs. Currently, image-text pair datasets (Schuhmann et al., 2022) and visual QA (Hudson and Manning, 2019; Mishra et al., 2019) data are used for training MLLMs. Although these datasets are usually larger than typical datasets in computer vision, they are still far less abundant than the text-only data used for training LLMs in terms of quantity. Insufficient data could potentially lead to problematic cross-modal alignment, resulting in hallucinations (Wang et al., 2023c; Sun et al., 2023).
@@ 3.1.2. Quality
  • Detailed descriptions (open question) The impact of the level of detail in textual descriptions on this matter remains an open question. As discussed in Sec. 2.2, the texts in pre-training data, such as LAION (Schuhmann et al., 2022), usually describe the salient objects’ overall content. While the texts in the instructing tuning stage, such as LLaVA-150k (Liu et al., 2023c), consist of more detailed descriptions. This LLaVA-150k dataset is generated by GPT-4 based on objects recognized by vision models. One recent work (Chen et al., 2023c) argues that within the training data, detailed descriptions related to object position, attributes, and non-salient objects are usually absent. This property results in incomplete cross-modal alignment and deprives the model of grounding ability (Zhang et al., 2023d; Lai et al., 2023). However, another work (Yue et al., 2024) hypothesizes that the text descriptions in the instruction tuning data contain too much details, exceeding the perception limit of MLLMs. When trained with such detailed data, in an attempt to fit the detail level and length distribution of ground truth captions, the model may risk expressing details that it cannot discern from the image, and therefore exhibit hallucinations. The detail level of the training data remains an open question.
@@ 3.2. Model
  • Weak vision model. As mentioned in related works (Rohrbach et al., 2018; Guan et al., 2023; Wang et al., 2023c), a primary potential reason for hallucination is a weak vision model, which can lead to misclassification or misinterpretation of visual concepts. Even the most powerful vision model may still experience information loss during the encoding process. Weak vision model implies weak perception, which fundamentally undermines the multimodal understanding.
@@ 3.3. Training
The training objective of MLLMs is basically the same as LLMs, i.e, auto-regressive next token prediction loss. This loss is straightforward yet effective and easy to scale up, showing promising performance in language modeling. However, some studies in the field of MLLMs have suggested that the next-token prediction loss might not be suitable for learning visual content due to its complex spatial structure (Ben-Kish et al., 2023; Chen et al., 2023c). Additionally, the loss optimizes at the token level, while lacking supervision at the sequence level (Ben-Kish et al., 2023). Another perspective is that, unlike training LLMs, the RLHF stage is absent in training procedure of MLLMs (Yu et al., 2023c; Sun et al., 2023), becoming a potential cause of hallucination.
@@ 3.4. Inference
As for inference, some works also argues a potential issue in the auto-regressive generation. During generation, as the sequence length grows, the self-attention will focus more on the previously generated text tokens, i.e., the attention on the visual content is diluted (Wang et al., 2023c, e; Huang et al., 2023b; Wang et al., 2023d). Through visualizing the attention map during generation (Huang et al., 2023b; Wang et al., 2023e), it can be observed that the generated content focuses more on previous special tokens, such as punctuation, rather than visual content tokens. The issue of ’losing attention’ would also lead to the model’s output response being irrelevant to the visual content.
@@ 4. Hallucination Metrics and Benchmarks
MME (Yin et al., 2023a) arXiv’23 Jun MSCOCO (Lin et al., 2014) 1457 Dis Acc/Score
@@
M-HalDetect (Gunjal et al., 2023) arXiv’23 Aug. MSCOCO (Lin et al., 2014) 4,000 Dis Reward Model Score
@@
NOPE (Lovenia et al., 2023) arXiv’23 Oct. Open-Images (Kuznetsova et al., 2020) 36,000 Dis Acc/METEOR (Banerjee and Lavie, 2005)
HaELM (Wang et al., 2023e) arXiv’23 Oct. MSCOCO (Lin et al., 2014) 5,000 Gen LLM Assessment Not Explicitly Stated
FaithScore (Jing et al., 2023) arXiv’23 Nov. MSCOCO (Lin et al., 2014) 2,000 Gen FaithScore Obj. Counting
@@
AMBER (Wang et al., 2023c) arXiv’23 Nov. Web 15,202 Dis & Gen AMBER Score
RAH-Bench (Chen et al., 2023c) arXiv’23 Nov. MSCOCO (Lin et al., 2014) 3,000 Dis False Positive Rate
@@
FGHE (Wang et al., 2023a) arXiv’23 Dec. MSCOCO (Lin et al., 2014) 200 Dis Acc/P/R/F Obj. Behavior
MOCHa (Ben-Kish et al., 2023) arXiv’23 Dec. Synthetic 2,000 Gen OpenCHAIR (Ben-Kish et al., 2023)
@@
VQAv2-IDK (Cha et al., 2024) arXiv’24 Feb. VQAv2 (Goyal et al., 2017) 6,624 Dis Acc IK (Cha et al., 2024)
MHaluBench (Chen et al., 2024a) arXiv’24 Feb. MSCOCO (Lin et al., 2014) 1,860 Gen Acc/P/R/F T2I
VHTest (Huang et al., 2024) arXiv’24 Feb. MSCOCO (Lin et al., 2014) 1,200 Dis & Gen Acc
Hal-Eavl (Jiang et al., 2024) arXiv’24 Feb. MSCOCO (Lin et al., 2014) & LAION (Schuhmann et al., 2022) 10,000 Dis & Gen Acc/P/R/F & LLM Assessment Obj. Event
@@
HaELM (Wang et al., 2023e) Most LLM-based evaluation benchmarks employ advanced ChatGPT or GPT-4 models to assess the quality of the MLLM response. In contrast, the work of Hallucination Evaluation based on Large Language Models (HaELM) proposes to train a specialized LLM for hallucination detection. It collects a set of hallucination data generated by a wide range of MLLMs, simulates data using ChatGPT, and trains an LLM based on LLaMA (Touvron et al., 2023). After that, the HaELM model becomes proficient in hallucination evaluation, leveraging reference descriptions of images as the basis of assessment.
@@
AMBER (Wang et al., 2023c) Upon the application and evaluation of MLLMs, the tasks can be roughly divided into generative tasks and discriminative tasks. For generative tasks, this paper argues that most existing works rely on additional LLMs, suffering from computational cost. As for discriminative tasks, the most popular evaluation suite is POPE (Li et al., 2023a). However, POPE lacks fine-grained hallucination types such as attributes and relations. AMBER (An LLM-free Multi-dimensional Benchmark) is proposed to support the evaluation of generative tasks and discriminative tasks, including object existence hallucination, attribute hallucination, and relation hallucination. It further combines the CHAIR (Rohrbach et al., 2018) metric in generative tasks and F1 in discriminative tasks to form the AMBER Score as follows:
@@
RAH-Bench (Chen et al., 2023c) Relation-Associated Hallucination Benchmark (RAH-Bench) can be regarded as an upgraded version of POPE, containing 3,000 yes-or-no questions with their corresponding images. Different from POPE, RAH-Bench further divides the negative questions into three subsets. Each subset contains 500 questions with misleading statements in the different aspects, including: 1) categorical hallucination, 2) attribute hallucination, 3) relation hallucination.
@@
FGHE (Wang et al., 2023a) Fine-Grained Object Hallucination Evaluation (FGHE) follows a binary classification approach similar to POPE to evaluate MLLMs. However, unlike POPE, FGHE requires a different set of binary questions to measure fine-grained hallucination. The FGHE dataset consists of 50 images and 200 binary questions divided into three categories: (a) multiple-object questions, which verify the relationships between multiple objects in the image; (b) attribute questions, which verify attributes of objects in the image; and (c) behavior questions, which verify behaviors or objects in the image. The questions are manually defined by human annotators on a subset of 50 images from the validation set of the MSCOCO dataset. Similar to POPE, the Accuracy, Precision, Recall, and F1 score are employed as the evaluation metrics.
@@
Hal-Eval (Jiang et al., 2024) The work of Hal-Eval (Jiang et al., 2024) identifies another type of object hallucination: event hallucination. This type of hallucination fabricates a fictional target and constructing an entire narrative around it, including its attributes, relationships, and actions. This effort further completes the definition of hallucination types. In addition, this work proposes an evaluation benchmark, which encompasses both discriminative and generative evaluation methods. This is achieved by collecting two evaluation subsets, each tailored to the discriminative and generative evaluation methods, respectively.
@@
MHaluBench (Chen et al., 2024a) This benchmark does not aim to evaluate the MLLMs themselves. Instead, it is intentionally designed to evaluate the hallucination detection tools of MLLMs, i.e., judge whether a tool can successfully detect the hallucination produced by an MLLM. Thus, the benchmark consists of hallucinatory examples. Specifically, the benchmark unifies image-to-text tasks and the text-to-image tasks into one evaluation suite: cross-modal consistency checking. The hallucinatory examples are generated using leading MLLMs and image generation models, such as LLaVA (Liu et al., 2023c), MiniGPT-4 (Zhu et al., 2023a), DALL-E2 (Ramesh et al., 2022), and DALL-E3 (Betker et al., 2023). During evaluation, the benchmark can be used to compare different hallucination detection methods based on their performance. So far, there are not many dedicated hallucination detection methods. This work serves as a basis for this direction.
@@
LLaVA-1.5 (Liu et al., 2023c) 7B - - - - - - - - 74.4 82.9 48.9 34.2
CogVLM (Wang et al., 2023b) 7B 195.00 165.00 103.33 160.00 - - - - 80 86.1 - -

### v2

@@ 2.3. Hallucinations in Multimodal Large Language Models
It’s worth noting that some literature may categorize objects counting, objects event, etc., as independent hallucination categories. In this work, we classify them under the attribute category. Some works also discuss ‘verb’ hallucination (Wang et al., 2024e), which might falls under the domain of either attribute or relation. Thus, we do not discuss it independently. The definition of hallucination types aligns well with the domain of compositional generalization (Yuksekgonul et al., 2022; Ma et al., 2023) of VLMs, which investigates visio-linguistic generalization and reasoning abilities.
@@ 3.1.1. Quantity
Deep learning models are data-hungry, especially large models like MLLMs. The amount of data plays an important role in building robust and reliable MLLMs. Currently, image-text pair datasets (Schuhmann et al., 2022) and visual QA (Hudson and Manning, 2019; Mishra et al., 2019) data are used for training MLLMs. Although these datasets are usually larger than typical datasets in computer vision, they are still far less abundant than the text-only data used for training LLMs in terms of quantity. Insufficient data could potentially lead to problematic cross-modal alignment, resulting in hallucinations (Wang et al., 2023b; Sun et al., 2023).
@@ 3.1.2. Quality
  • Detailed descriptions (open question) The impact of the level of detail in textual descriptions on this matter remains an open question. As discussed in Sec. 2.2, the texts in pre-training data, such as LAION (Schuhmann et al., 2022), usually describe the salient objects’ overall content. While the texts in the instructing tuning stage, such as LLaVA-150k (Liu et al., 2023c), consist of more detailed descriptions. This LLaVA-150k dataset is generated by GPT-4 based on objects recognized by vision models. One recent work (Chen et al., 2023b) argues that within the training data, detailed descriptions related to object position, attributes, and non-salient objects are usually absent. This property results in incomplete cross-modal alignment and deprives the model of grounding ability (Zhang et al., 2023d; Lai et al., 2023). However, another work (Yue et al., 2024) hypothesizes that the text descriptions in the instruction tuning data contain too much details, exceeding the perception limit of MLLMs. When trained with such detailed data, in an attempt to fit the detail level and length distribution of ground truth captions, the model may risk expressing details that it cannot discern from the image, and therefore exhibit hallucinations. The detail level of the training data remains an open question.
@@ 3.2. Model
  • Weak vision model. As mentioned in related works (Rohrbach et al., 2018; Guan et al., 2023; Wang et al., 2023b), a primary potential reason for hallucination is a weak vision model, which can lead to misclassification or misinterpretation of visual concepts. Even the most powerful vision model may still experience information loss during the encoding process. Weak vision model implies weak perception, which fundamentally undermines the multimodal understanding.
@@ 3.3. Training
The training objective of MLLMs is basically the same as LLMs, i.e, auto-regressive next token prediction loss. This loss is straightforward yet effective and easy to scale up, showing promising performance in language modeling. However, some studies in the field of MLLMs have suggested that the next-token prediction loss might not be suitable for learning visual content due to its complex spatial structure (Ben-Kish et al., 2023; Chen et al., 2023b). Additionally, the loss optimizes at the token level, while lacking supervision at the sequence level (Ben-Kish et al., 2023). Another perspective is that, unlike training LLMs, the RLHF stage is absent in training procedure of MLLMs (Yu et al., 2023c; Sun et al., 2023), becoming a potential cause of hallucination.
@@ 3.4. Inference
3.4.1. Visual Attention Deficiency
As for inference, some works also argues a potential issue in the auto-regressive generation. During generation, as the sequence length grows, the self-attention will focus more on the previously generated text tokens, i.e., the attention on the visual content is diluted (Wang et al., 2023b, d; Huang et al., 2023b; Wang et al., 2023c; Favero et al., 2024). Through visualizing the attention map during generation (Huang et al., 2023b; Wang et al., 2023d), it can be observed that the generated content focuses more on previous special tokens, such as punctuation, rather than visual content tokens. The issue of ’losing attention’ would also lead to the model’s output response being irrelevant to the visual content.
3.4.2. Trap Visual Tokens
Even within the group of visual tokens, each token may play a different role. The work of AvisC (Woo et al., 2024b) analyzes that there are a few image tokens receiving excessive attention, while tokens receiving lower attention weights often hold essential information for identifying nuanced object details. This imbalanced attention distribution among visual tokens can lead to hallucinatory responses in tasks requiring fine-grained understanding of visual objects. More recently, the work of VTI (Liu et al., 2025) finds that vision encoder may generate some tokens that are very sensitive to image noise. Such noise can produce outliers in the representation space, resulting in large variability in some of the features, hence hallucinations.
Similar findings have also been discussed in some recent work (Che et al., 2025; Chen et al., 2025b), i.e., a small portion of visual tokens can severely affect the final results and induce hallucination.
@@ 4. Hallucination Metrics and Benchmarks
MME (Fu et al., 2023) arXiv’23 Jun MSCOCO (Lin et al., 2014) 1457 Dis Acc/Score
MMBench (Liu et al., 2024a) ECCV’24 Not Specified 3217 Dis Acc Reasoning
@@
M-HalDetect (Gunjal et al., 2024) AAAI’24 MSCOCO (Lin et al., 2014) 4,000 Dis Reward Model Score
@@
NOPE (Lovenia et al., 2023) ACL-W’24 Open-Images (Kuznetsova et al., 2020) 36,000 Dis Acc/METEOR (Banerjee and Lavie, 2005)
HaELM (Wang et al., 2023d) arXiv’23 Oct. MSCOCO (Lin et al., 2014) 5,000 Gen LLM Assessment Not Explicitly Stated
FaithScore (Jing et al., 2023) EMNLP’24 MSCOCO (Lin et al., 2014) 2,000 Gen FaithScore Obj. Counting
@@
AMBER (Wang et al., 2023b) arXiv’23 Nov. Web 15,202 Dis & Gen AMBER Score
RAH-Bench (Chen et al., 2023b) arXiv’23 Nov. MSCOCO (Lin et al., 2014) 3,000 Dis False Positive Rate
@@
FGHE (Wang et al., 2024d) MMM’24 MSCOCO (Lin et al., 2014) 200 Dis Acc/P/R/F Obj. Behavior
MOCHa (Ben-Kish et al., 2023) EMNLP’24 Synthetic 2,000 Gen OpenCHAIR (Ben-Kish et al., 2023)
@@
VQAv2-IDK (Cha et al., 2024) ICASSP’24 VQAv2 (Goyal et al., 2017) 6,624 Dis Acc IK (Cha et al., 2024)
MHaluBench (Chen et al., 2024c) ACL’24 MSCOCO (Lin et al., 2014) 1,860 Gen Acc/P/R/F T2I
VHTest (Huang et al., 2024) ACL’24 MSCOCO (Lin et al., 2014) 1,200 Dis & Gen Acc
Hal-Eavl (Jiang et al., 2024a) MM’24 MSCOCO (Lin et al., 2014) & LAION (Schuhmann et al., 2022) 10,000 Dis & Gen Acc/P/R/F & LLM Assessment Obj. Event
PhD (Liu et al., 2024b) arXiv’24 Mar. TDIUC (Kafle and Kanan, 2017) & AIGC 102,564 Dis PhD Index Sentiment
THRONE (Kaul et al., 2024) CVPR’24 MSCOCO (Lin et al., 2014) 5,000, Gen P/R/F
BEAF (Ye-Bin et al., 2024) ECCV’24 MSCOCO (Lin et al., 2014) 26,118 Dis TU/IG/SB/ID
ROPE (Chen et al., 2024b) NeurIPS’24 MSCOCO (Lin et al., 2014) & ADE20k (Zhou et al., 2019) 5,000 Dis Acc Multi Obj.
LongHalQA (Qiu et al., 2024) arXiv’24 Oct. VisualGenome (Krishna et al., 2017) & Objects365 (Shao et al., 2019) 6,485 Dis & Gen Acc Obj. Counting
Reefknot (Zheng et al., 2024a) arXiv’24 Dec. VisualGenome (Krishna et al., 2017) 21,880 Dis & Gen R-score
@@
HaELM (Wang et al., 2023d) Most LLM-based evaluation benchmarks employ advanced ChatGPT or GPT-4 models to assess the quality of the MLLM response. In contrast, the work of Hallucination Evaluation based on Large Language Models (HaELM) proposes to train a specialized LLM for hallucination detection. It collects a set of hallucination data generated by a wide range of MLLMs, simulates data using ChatGPT, and trains an LLM based on LLaMA (Touvron et al., 2023). After that, the HaELM model becomes proficient in hallucination evaluation, leveraging reference descriptions of images as the basis of assessment.
@@
AMBER (Wang et al., 2023b) Upon the application and evaluation of MLLMs, the tasks can be roughly divided into generative tasks and discriminative tasks. For generative tasks, this paper argues that most existing works rely on additional LLMs, suffering from computational cost. As for discriminative tasks, the most popular evaluation suite is POPE (Li et al., 2023a). However, POPE lacks fine-grained hallucination types such as attributes and relations. AMBER (An LLM-free Multi-dimensional Benchmark) is proposed to support the evaluation of generative tasks and discriminative tasks, including object existence hallucination, attribute hallucination, and relation hallucination. It further combines the CHAIR (Rohrbach et al., 2018) metric in generative tasks and F1 in discriminative tasks to form the AMBER Score as follows:
@@
RAH-Bench (Chen et al., 2023b) Relation-Associated Hallucination Benchmark (RAH-Bench) can be regarded as an upgraded version of POPE, containing 3,000 yes-or-no questions with their corresponding images. Different from POPE, RAH-Bench further divides the negative questions into three subsets. Each subset contains 500 questions with misleading statements in the different aspects, including: 1) categorical hallucination, 2) attribute hallucination, 3) relation hallucination.
@@
FGHE (Wang et al., 2024d) Fine-Grained Object Hallucination Evaluation (FGHE) follows a binary classification approach similar to POPE to evaluate MLLMs. However, unlike POPE, FGHE requires a different set of binary questions to measure fine-grained hallucination. The FGHE dataset consists of 50 images and 200 binary questions divided into three categories: (a) multiple-object questions, which verify the relationships between multiple objects in the image; (b) attribute questions, which verify attributes of objects in the image; and (c) behavior questions, which verify behaviors or objects in the image. The questions are manually defined by human annotators on a subset of 50 images from the validation set of the MSCOCO dataset. Similar to POPE, the Accuracy, Precision, Recall, and F1 score are employed as the evaluation metrics.
@@
Hal-Eval (Jiang et al., 2024a) The work of Hal-Eval (Jiang et al., 2024a) identifies another type of object hallucination: event hallucination. This type of hallucination fabricates a fictional target and constructing an entire narrative around it, including its attributes, relationships, and actions. This effort further completes the definition of hallucination types. In addition, this work proposes an evaluation benchmark, which encompasses both discriminative and generative evaluation methods. This is achieved by collecting two evaluation subsets, each tailored to the discriminative and generative evaluation methods, respectively.
@@
MHaluBench (Chen et al., 2024c) This benchmark does not aim to evaluate the MLLMs themselves. Instead, it is intentionally designed to evaluate the hallucination detection tools of MLLMs, i.e., judge whether a tool can successfully detect the hallucination produced by an MLLM. Thus, the benchmark consists of hallucinatory examples. Specifically, the benchmark unifies image-to-text tasks and the text-to-image tasks into one evaluation suite: cross-modal consistency checking. The hallucinatory examples are generated using leading MLLMs and image generation models, such as LLaVA (Liu et al., 2023c), MiniGPT-4 (Zhu et al., 2023a), DALL-E2 (Ramesh et al., 2022), and DALL-E3 (Betker et al., 2023). During evaluation, the benchmark can be used to compare different hallucination detection methods based on their performance. So far, there are not many dedicated hallucination detection methods. This work serves as a basis for this direction.
@@
THRONE (Kaul et al., 2024) THRONE categorizes hallucinations in generative and discriminative tasks as Type I and Type II hallucination. It argues that most existing benchmarks focus on Type II and result does not translate into Type I. Regarding evaluating Type I, the main drawback of existing works is the reliance on close-source GPT model. Thus, this works proposes an evaluation framework based on MSCOCO dataset, which 1) evaluates object hallucination in free-form output; 2) utilizes open-source LLM instead of close-source ones.
BEAF (Ye-Bin et al., 2024) BEAF curates a new evaluation dataset, called the BEfore-AFter hallucination dataset (BEAF), and introduces new metrics: True Understanding(TU), IGnorance (IG), StuBbornness (SB), and InDecision (ID). Different from prior works that focus only on constructing questions and answers, the key idea of this benchmark is to manipulate visual scene information by image editing models and to design the metrics based on scene changes. This allows clearly assessment on whether VLMs correctly understand a given scene by observing the ability to perceive changes.
@@
LLaVA-1.5 (Liu et al., 2023c) 7B - - - - 83.44 82.08 79.26 - 74.4 82.9 48.9 34.2
CogVLM (Wang et al., 2023a) 7B 195.00 165.00 103.33 160.00 - - - - 80 86.1 - -
Intern-VL2 (Chen et al., 2024d) 7B - - - - 82.60 81.64 80.70 - - - - -
DeepSeek-VL2 (Wu et al., 2024a) 7B - - - - 87.64 85.86 85.55 - - - - -
Qwen-VL2 (Wang et al., 2024a) 7B - - - - 88.99 87.68 86.02 - - - - -
LLaVA-One-Vision (Li et al., 2024) 7B - - - - 88.51 87.33 86.27 - - - - -
OPERA(LLaVA-1.5) (Huang et al., 2023b) 7B 180.67 133.33 123.33 155.00 89.95 86.88 81.77 - - - - -
VCD(LLaVA-1.5) (Leng et al., 2023) 7B 186.67 125.56 128.89 139.45 87.15 83.94 79.47 - - - - -
M3ID(LLaVA-1.5) (Favero et al., 2024) 7B 186.67 128.33 131.67 151.67 87.52 84.95 80.22 - - - - -
RITUAL(LLaVA-1.5) (Woo et al., 2024a) 7B 187.50 139.58 125 164.17 88.81 86.17 80.54
DeGF(LLaVA-1.5) (Zhang et al., 2025b) 7B 188.33 150.00 133.89 172.22 88.74 86.28 81.94 - - - - -
SID(LLaVA-1.5) (Huo et al., 2024) 7B 190.00 148.33 128.33 175.00 89.00 85.01 81.28 - - - - -
CCA(LLaVA-1.5) (Xing et al., 2024a) 7B 190.00 148.33 128.33 175.00 89.05 86.02 83.82 - - - - -
DAC(LLaVA-1.5) (Zhang et al., 2025a) 7B 195.00 158.33 133.33 170.00 90.60 89.10 84.42 - - - - -